# Fermion liquids as quantum Hall liquids in phase space: A unified approach for anomalies and responses

### v1

@@ II.1.2 First order solution in $\theta$
To illustrate the systematic procedures in determining the SW map, we discuss in detail the first order SW map here. It starts by deducing $\Lambda_{\lambda}^{(1)}$ since Eq. (15) only involves $\Lambda$ and $\lambda$. By expanding Eq. (15) up to the first order in $\theta$, we obtain
@@
For the gauge potential, performing a similar series expansion for Eq. (16) and retaining terms up to the first order in $\theta$ gives:
@@
$A_{\mu,3D}^{(1)}=-\frac{1}{2}\partial_{\beta}\left(\theta^{\alpha\beta}A_{\alpha}A_{\mu}\right)-\frac{\hbar}{2}\delta_{\mu,0}\epsilon^{\alpha\beta\nu}A_{\alpha}\partial_{\beta}A_{\nu}.$ (30)
Note that in general, setting $\hat{A}_{\mu}=0$ does not exclude the possibility of the corresponding commutative gauge field $A_{\mu}$ being non-zero under the SW map. When we go to the phase space, this allows us to set the noncommutative fields $\hat{A}_{p_{i}}=\hat{A}_{x_{i}}=0$ for simplicity.
Lastly, we consider the first order expansion for the matter field using Eq. (II.1.1):
@@
Since $\psi^{(1)}$ appears on both sides, $\psi^{(1)}$ must be a product between an expression and $\psi$. After the action of $\delta_{\lambda}$, the expression should cancel the remaining two terms on the right. This suggests a possible inhomogeneous solution JMSSW2001 ; Ulker2008
@@ III Action for phase-space fermion liquid in arbitrary dimensions
Assuming the usual non-commutativity relations (Eq. (7)) for the position $X_{i}$ and momentum $P_{i}$ variables, we propose the following Euclidean action for the incompressible liquid in phase space,
@@
where the ellipses include terms of quartic and higher order in the fermions. In writing the above action, we have already traded the non-commutativity in the coordinates for the Moyal-Weyl star product. The (classical) phase space measure is
@@
In a fully translation invariant system the fermion liquid has a boundary in the momentum directions corresponding to the Fermi surface, and no boundary in real space because periodic boundary conditions need to be imposed in real-space directions to maintain translation invariance. In reality, on the other hand, there is a real space boundary which breaks translation invariance. The advantage of our phase-space formalism is that boundaries in real-space and momentum-space, which are both parts of phase space boundaries, can be treated on the same footing. In order to contain the liquid, there must be a confining potential in all directions in $\hat{A}_{0}(x_{i},p_{i},\tau)$, which is nothing but the non-interacting Hamiltonian for the fermion liquid:
@@
where $\epsilon(\textbf{p})$ is the band dispersion and $\hat{V}$ is the external potential (which could in principle be momentum dependent). So in our phase space formulation the kinetic (band) energy is nothing but a confining potential in momentum space, which is treated on equal footing with real space potential. For simplicity, we set $\hat{A}_{p}=\hat{A}_{x}=0$.
@@ IV Emergence of the Chern-Simons term in 1D
Hereafter, we assume $\Psi$ and $\psi$ are time-independent. As $\bar{\psi}\psi\approx\bar{\rho}\sim\hbar^{-1}$, the term written explicitly in Eq. (39) is of the order $\mathcal{O}(1/\hbar)$.
Next, we consider the first order $\theta$ expansion for $\hat{A}_{\mu}$ in Eq. (28). The zeroth component is given by
@@
The existence of boundary terms begets the question of gauge invariance of the effective actions. Of course, the starting action Eq. (38) is explicitly gauge invariant under a U(1) gauge transformation in phase space. The full SW map preserves this gauge invariance. The constraint $\partial_{\tau}\Psi=0$ is a gauge choice, which explains the apparent lack of gauge invariance at order $\mathcal{O}(1/\hbar)$ in $S_{-1}$. Additionally, $S_{0,B}$ is exactly cancelled by a term that one gets upon relaxing incompressibility,
@@ V.1.1 Contribution from first order expansion for $\hat{A}_{0}$
This is similar to having two decoupled Landau levels in $(x,p_{x})$ and $(y,p_{y})$ Yang-PRA2020 . Note that $S_{-1}$ is not a genuine Chern-Simons term in the $5$-dimensional spacetime, but the two 3-dimensional subspaces $(\tau,x,p_{x})$ and $(\tau,y,p_{y})$. $S_{-1,B}$ collects the boundary terms from the two lower-dimensional CS forms.
The CS 3-forms in $S_{-1}$ are ‘non-topological’ in the sense that their coefficients are not quantized in 4+1-D phase space. However, this does not mean that their forms are arbitrary. In fact, $S_{-1}$ is invariant under any canonical transformations in phase space. Under a canonical transformation
@@
The same holds true for the remaining terms in the CS 3-form as well. Thus, the total contribution to the physical responses from the non-topological terms are invariant under arbitrary canonical transformations in phase space.
@@ V.1.2 Contribution from second order expansion for $\hat{A}_{0}$
The topological CS term describes several contributions to responses in the 2D EFL Qi2015 ; ETS , especially the unquantized part of the anomalous Hall response AHEReview in the presence of non-zero Berry curvature $F_{p_{x},p_{y}}$ [^2] The momentum-space gauge fields $A_{p_{x}},A_{p_{y}}$ correspond to the Berry connection Qi2015 .. Additionally, if we view the current as being generated by Bloch wavepackets $\phi_{\textbf{R},\textbf{K}}(x,y)$, we also obtain the response due to mixed space Berry curvatures DiXiao2010 ; Sundaram1999 ; Randeria2024 , such as
@@ V.2 Semiclassical terms in the response
An upshot of our treatment of the phase space quantum Hall liquid is that in addition to the topological CS term in 4+1-D phase space proposed by ETS, we also get the lower dimensional (and non-topological) CS forms (Eq. (50)). The CS 3-forms ($\mathrm{CS}_{3}$) generate the group velocity term in the current,
@@
As noted earlier, the $p$-dependent part of $A_{0}$ is the dispersion $\epsilon(p)$ and thus $\partial_{p_{i}}A_{0}=\partial_{p_{i}}\epsilon$ is the group velocity. Adopting the electronic wavepacket point of view, we also get its momentum-space equivalent
$J_{p_{i}}^{\mathrm{CS}_{3}}\propto\frac{dK_{i}}{dt}=-\partial_{x_{i}}A_{0}$ (58)
@@ VI Fermion liquids in 3D and beyond
Figure 2: Generation of different CS actions in the 6D phase space with imaginary time, and the corresponding subspaces with smaller odd dimensions under the SW map. Note that there is only a single imaginary time, $\tau$.
For the 3D fermion liquids, similar to the previous cases, the lowest order expansion consists of several terms (Fig. 2)
@@
$S_{-1}[A]=\frac{-i\Omega_{z,p_{z}}}{6(2\pi)^{3}\hbar}\int d^{5}x\ A\wedge dA\wedge dA\\ +(x\rightarrow y\rightarrow z)+(x\rightarrow z\rightarrow y)+S_{-1,B}+\mathcal{O(\delta\rho)}$ (61)
@@
$S_{0}-S_{0,B}=\frac{i}{(2\pi)^{3}4!}\int d^{7}x\ A\wedge dA\wedge dA\wedge dA$ (62)
@@
Again, like the 2D scenario, we obtain the non-topological terms $S_{-1}$ and $S_{-2}$ which describe physically relevant responses of the fermion liquids. $S_{-2}$ contains three sets of CS 3-forms for each pair of conjugate variables $(x_{i},p_{i})$ and generates the semiclassical responses Eqs. (57) and (58). The set of CS 5-forms $S_{-1}$ is responsible for generating the anomalous Hall responses in all three spatial currents $J_{x},J_{y},$ and $J_{z}$ when the Fermi surface encloses non-zero Berry flux
@@
The non-quantized part of the anomalous Hall response in the equilibrium current is yet another important contribution that cannot be obtained from the topological CS term (Eq. 62) that yields the ETS anomaly. Thus, our approach unifies all possible responses of the fermion liquid to real- or momentum-space gauge fields.
@@ VII Concluding Remarks
We have shown that fermion liquids with Fermi surfaces satisfying Luttinger sum rule can be viewed as incompressible liquids in phase-space, which are topologically non-trivial in the way similar to quantum Hall liquids. Taking into consideration the non-commutativity of phase space, we employed the Seiberg-Witten map to derive the effective action governing their responses to external probing gauge fields. In addition to the topological term that captures the ETS anomaly, the phase-space formalism also yields lower-order terms that reproduce semiclassical and other non-quantized responses in 2D and 3D. From the viewpoint of non-commutative field theory, our work also presents an important technical advancement. It has been a longstanding conjecture Picariello2002 that CS forms arise at all orders in the SW map. We prove that this is indeed true, up to order 3. In Appendix C, we sketch a proof by induction for the higher-order terms.
@@
Our work unifies the description of Fermi and non-Fermi liquids, as they share the same bulk (in phase space) topological properties. Their difference, on the other hand, lies at the phase space boundary, including the Fermi surface. This is similar to what happens in quantum Hall liquids with a specific bulk topological order, where edge physics can still be different due to non-universal physics like edge reconstruction. As a concrete example, the edge of a $\nu=1$ quantum Hall liquid is a Fermi liquid without reconstruction, but becomes a Luttinger liquid with reconstruction. Our work further unifies the description of Fermi surface (or momentum space boundary) with real space boundary, as they are both parts of phase space boundary. Strictly speaking, Fermi surfaces are not sharp in the presence of disorder and/or real space boundary that break translation symmetry; one can only talk about phase space boundary in such cases. This suggests the possibility of probing Fermi surface physics at the real space boundary of the system, in a way similar to edge physics of quantum Hall liquids.
@@
In a way our work also unifies the description of metals and (band or Mott) insulators, which are normally considered the exact opposite of each other. The similarity between them, in the context of the present paper, lies in their incompressibility, one in phase space and the other in real space. In fact insulators are also incompressible in phase space, except their boundaries are confined to real space (in other words, they have no Fermi surface!). From this viewpoint the difference between metal and insulator lies in the (phase-space) boundary, with insulators being simpler as their phase space boundaries are orthogonal to the momentum space.
@@
This line of work will likely lead to new insights into high-dimensional bosonization (of Fermi surface), the formulation of which has not been entirely satisfactory. A promising new formulation has appeared recently Son-PRR2022 . Our perspective, namely treating the Fermi surface as the boundary of an incompressible liquid in phase space, is complementary to this approach.
@@
It is tempting to generalize our approach to study interacting quantum Hall liquid in phase space that are topologically distinct from those studied here, and explore its physical consequences. In particular, one can consider fractional version of the incompressible liquids with higher levels of CS term in their effective action. Such liquids satisfy a fractional version of the Luttinger sum rule, in a way similar to fractional quantum Hall liquids with a fractional bulk filling factor. Due to the quantization of the level of the CS term, the phase space density of such liquids must be a rational fraction of that dictated by the Luttinger sum rule, the simplest version of which is $1/m$ with $m$ being an integer. This corresponds to the Laughlin sequence in fractional quantum Hall effect. It is not immediately clear, however, whether $m$ has to be odd for $d\geq 2$. We leave these issues to be addressed in future work.
@@ Acknowledgements.
The authors acknowledge helpful conversations with T. Senthil. J.P. acknowledges the support from Florida State University through the FSU Quantum Postdoctoral Fellowship. This research is supported by the National Science Foundation Grant No. DMR-2315954. Most of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556, and the State of Florida.
@@ Appendix B Chern-Simons term in the third order SW map
In this Appendix, we show that the diagonal term in the third-order SW map for the gauge potential, $A_{\mu}^{(3,3)}$ in non-commutative phase space can be written as a sum of a total derivative and a term of the Chern-Simons form. For the case of 6+1-D phase-space, we obtain the topological CS term. From the recursive relation for diagonal terms Fidanza , we have
@@
with the understanding that the $\epsilon^{\mu\nu\sigma\rho}$ antisymmetrizes non-zero indices in groups of four. For example, in 6+1-D, it will antisymmetrize the three groups
@@
when the Greek indices take non-zero values $(1,\ldots,6)$. For the first term in Eq. (98), we verified the identity

### v2

@@ II.1.2 First order solution in $\theta$
To illustrate the systematic procedures in determining the SW map, we discuss in detail the first-order SW map here. It starts by deducing $\Lambda_{\lambda}^{(1)}$ since Eq. (15) only involves $\Lambda$ and $\lambda$. By expanding Eq. (15) up to the first order in $\theta$, we obtain
@@
For the gauge potential, performing a similar series expansion for Eq. (16) and retaining terms up to the first order in $\theta$ gives
@@
$A_{\mu,\text{3D}}^{(1)}=-\frac{1}{2}\partial_{\beta}\left(\theta^{\alpha\beta}A_{\alpha}A_{\mu}\right)-\frac{\hbar}{2}\delta_{\mu,0}\epsilon^{\alpha\beta\nu}A_{\alpha}\partial_{\beta}A_{\nu}.$ (30)
Note that in general, setting $\hat{A}_{\mu}=0$ does not exclude the possibility of the corresponding commutative gauge field $A_{\mu}$ being nonzero under the SW map. When we go to the phase space, this allows us to set the noncommutative fields $\hat{A}_{p_{i}}=\hat{A}_{x_{i}}=0$ for simplicity.
Lastly, we consider the first-order expansion for the matter field using Eq. (II.1.1):
@@
Since $\psi^{(1)}$ appears on both sides, $\psi^{(1)}$ must be a product between an expression and $\psi$. After the action of $\delta_{\lambda}$, the expression should cancel the remaining two terms on the right. This suggests a possible inhomogeneous solution JMSSW2001 ; Ulker2008 ,
@@ III Action for phase-space fermion liquid in arbitrary dimensions
Assuming the usual noncommutativity relations (Eq. (7)) for the position $X_{i}$ and momentum $P_{i}$ variables, we propose the following Euclidean action for the incompressible liquid in phase space,
@@
where the ellipses include terms of quartic and higher order in the fermions. In writing the above action, we have already traded the noncommutativity in the coordinates for the Moyal-Weyl star product. The (classical) phase space measure is
@@
In a fully translation-invariant system the fermion liquid has a boundary in the momentum directions corresponding to the Fermi surface, and no boundary in real space because periodic boundary conditions need to be imposed in real-space directions. In reality, on the other hand, there is a real-space boundary which breaks translation invariance. The advantage of our phase-space formalism is that boundaries in real space and momentum space, which are both parts of phase-space boundaries, can be treated on the same footing. In order to contain the liquid, there must be a confining potential in all directions in $\hat{A}_{0}(x_{i},p_{i},\tau)$, which is nothing but the noninteracting Hamiltonian for the fermion liquid:
@@
where $\epsilon(\textbf{p})$ is the band dispersion and $\hat{V}$ is the external potential (which could in principle be momentum dependent). So, in our phase-space formulation the kinetic (band) energy is nothing but a confining potential in momentum space, which is treated on equal footing with real-space potential. For simplicity, we set $\hat{A}_{p}=\hat{A}_{x}=0$.
@@ IV Emergence of the Chern-Simons term in 1D
Hereafter, we assume $\Psi$ and $\psi$ are time independent. As $\bar{\psi}\psi\approx\bar{\rho}\sim\hbar^{-1}$, the term written explicitly in Eq. (39) is of the order $\mathcal{O}(1/\hbar)$.
Next, we consider the first-order $\theta$ expansion for $\hat{A}_{\mu}$ in Eq. (28). The zeroth component is given by
@@
The existence of boundary terms begets the question of gauge invariance of the effective actions. Of course, the starting action Eq. (38) is explicitly gauge invariant under a U(1) gauge transformation in phase space. The full SW map preserves this gauge invariance. The constraint $\partial_{\tau}\Psi=0$ is a gauge choice, which explains the apparent lack of gauge invariance at order $\mathcal{O}(1/\hbar)$ in $S_{-1}$. Additionally, $S_{0,B}$ is exactly canceled by a term that one gets upon relaxing incompressibility,
@@ V.1.1 Contribution from first order expansion for $\hat{A}_{0}$
This is similar to having two decoupled Landau levels in $(x,p_{x})$ and $(y,p_{y})$ Yang-PRA2020 . Note that $S_{-1}$ is not a genuine Chern-Simons term in the five-dimensional spacetime, but the two three-dimensional subspaces $(\tau,x,p_{x})$ and $(\tau,y,p_{y})$. $S_{-1,B}$ collects the boundary terms from the two lower-dimensional CS forms.
The CS 3-forms in $S_{-1}$ are “nontopological” in the sense that their coefficients are not quantized in 4+1-D phase space. However, this does not mean that their forms are arbitrary. In fact, $S_{-1}$ is invariant under any canonical transformations in phase space. Under a canonical transformation
@@
The same holds true for the remaining terms in the CS 3-form as well. Thus, the total contribution to the physical responses from the nontopological terms are invariant under arbitrary canonical transformations in phase space.
@@ V.1.2 Contribution from second order expansion for $\hat{A}_{0}$
The topological CS term describes several contributions to responses in the 2D EFL Qi2015 ; ETS , especially the unquantized part of the anomalous Hall response AHEReview in the presence of nonzero Berry curvature $F_{p_{x},p_{y}}$ [^2] The momentum-space gauge fields $A_{p_{x}},A_{p_{y}}$ correspond to the Berry connection Qi2015 .. Additionally, if we view the current as being generated by Bloch wavepackets $\phi_{\textbf{R},\textbf{K}}(x,y)$, we also obtain the response due to mixed space Berry curvatures DiXiao2010 ; Sundaram1999 ; Randeria2024 , such as
@@ V.2 Semiclassical terms in the response
An upshot of our treatment of the phase-space quantum Hall liquid is that in addition to the topological CS term in 4+1-D phase space proposed by ETS, we also get the lower-dimensional (and nontopological) CS forms [Eq. (50)]. The CS 3-forms ($\mathrm{CS}_{3}$) generate the group velocity term in the current,
@@
As noted earlier, the $p$-dependent part of $A_{0}$ is the dispersion $\epsilon(p)$ and thus $\partial_{p_{i}}A_{0}=\partial_{p_{i}}\epsilon$ is the group velocity. Adopting the electronic wave-packet point of view, we also get its momentum-space equivalent
$J_{p_{i}}^{\mathrm{CS}_{3}}\propto\frac{dK_{i}}{dt}=-\partial_{x_{i}}A_{0},$ (58)
@@ VI Fermion liquids in 3D and beyond
Figure 2: Generation of different CS actions in the six-dimensional phase space with imaginary time, and the corresponding subspaces with smaller odd dimensions under the SW map. Note that there is only a single imaginary time, $\tau$.
For the three-dimensional fermion liquids, similar to the previous cases, the lowest order expansion consists of several terms (Fig. 2),
@@
$S_{-1}[A]=\frac{-i\Omega_{z,p_{z}}}{6(2\pi)^{3}\hbar}\left.\int d^{5}x\ \epsilon^{\mu\nu\sigma\rho\delta}A_{\mu}\partial_{\nu}A_{\sigma}\partial_{\rho}A_{\delta}\right|_{\begin{subarray}{c}~{}~{}(\mu,\nu,\sigma,\rho,\delta)\\ =(0,1,2,3,4)\end{subarray}}\\ +(x\rightarrow y\rightarrow z)+(x\rightarrow z\rightarrow y)+S_{-1,B}+\mathcal{O(\delta\rho)}$ (61)
@@
$S_{0}-S_{0,B}=\frac{i}{(2\pi)^{3}4!}\int A\wedge dA\wedge dA\wedge dA$ (62)
@@
Again, like the 2D scenario, we obtain the nontopological terms $S_{-1}$ and $S_{-2}$ which describe physically relevant responses of the fermion liquids. $S_{-2}$ contains three sets of CS 3-forms for each pair of conjugate variables $(x_{i},p_{i})$ and generates the semiclassical responses Eqs. (57) and (58). The set of CS 5-forms $S_{-1}$ is responsible for generating the anomalous Hall responses in all three spatial currents $J_{x},J_{y},$ and $J_{z}$ when the Fermi surface encloses nonzero Berry flux
@@
The nonquantized part of the anomalous Hall response in the equilibrium current is yet another important contribution that cannot be obtained from the topological CS term [Eq. 62] that yields the ETS anomaly. Thus, our approach unifies all possible responses of the fermion liquid to real- or momentum-space gauge fields.
@@ VII Concluding Remarks
We have shown that fermion liquids with Fermi surfaces satisfying the Luttinger sum rule can be viewed as incompressible liquids in phase space, which are topologically non-trivial in the way similar to quantum Hall liquids. Taking into consideration the noncommutativity of phase space, we employed the Seiberg-Witten map to derive the effective action governing their responses to external probing gauge fields. In addition to the topological term that captures the ETS anomaly, the phase-space formalism also yields lower-order terms that reproduce semiclassical and other nonquantized responses in 2D and 3D. From the viewpoint of noncommutative field theory, our work also presents an important technical advancement. It has been a longstanding conjecture Picariello2002 that CS forms arise at all orders in the SW map. We prove that this is indeed true, up to order 3. In Appendix C, we sketch a proof by induction for the higher-order terms.
@@
Our work unifies the description of Fermi and non-Fermi liquids, as they share the same bulk (in phase space) topological properties. Their difference, on the other hand, lies at the phase space boundary, including the Fermi surface. This is similar to what happens in quantum Hall liquids with a specific bulk topological order, where edge physics can still be different due to non-universal physics like edge reconstruction. As a concrete example, the edge of a $\nu=1$ quantum Hall liquid is a Fermi liquid without reconstruction, but becomes a Luttinger liquid with reconstruction. Our work further unifies the description of Fermi surface (or momentum-space boundary) with real-space boundary, as they are both parts of phase-space boundary. Strictly speaking, Fermi surfaces are not sharp in the presence of disorder and/or a real-space boundary. As both break translation symmetry, one can only talk about phase-space boundary in such cases. This suggests the possibility of probing Fermi surface physics at the real-space boundary of the system, in a way similar to edge physics of quantum Hall liquids.
@@
In a way our work also unifies the description of metals and (band or Mott) insulators, which are normally considered the exact opposite of each other. The similarity between them, in the context of the present paper, lies in their incompressibility, one in phase space and the other in real space. In fact, insulators are also incompressible in phase space, except their boundaries are confined to real space (in other words, they have no Fermi surface!). From this viewpoint the difference between metal and insulator lies in the (phase space) boundary, with insulators being simpler because their phase-space boundaries are orthogonal to the momentum space directions.
@@
This line of work will likely lead to new insights into high-dimensional bosonization (of the Fermi surface), the formulation of which has not been entirely satisfactory. A promising formulation has appeared recently in Ref.  Son-PRR2022 and subsequent works Huang_2024 ; Ye-2024 ; Ye-2024-2 . Our perspective, namely, treating the Fermi surface as the boundary of an incompressible liquid in phase space, is complementary to this approach.
@@
It is tempting to generalize our approach to study interacting quantum Hall liquid in phase spaces that are topologically distinct from those studied here, and explore its physical consequences. In particular, one can consider a fractional version of the incompressible liquids with higher levels of CS term in their effective action. Such liquids satisfy a fractional version of the Luttinger sum rule, in a way similar to fractional quantum Hall liquids with a fractional bulk filling factor. Due to the quantization of the level of the CS term, the phase space density of such liquids must be a rational fraction of that dictated by the Luttinger sum rule, the simplest version of which is $1/m$, with $m$ being an integer. This corresponds to the Laughlin sequence in fractional quantum Hall effect Agarwal-Karabali-2025 . It is not immediately clear, however, whether $m$ has to be odd for $d\geq 2$. We leave these issues to be addressed in future works.
@@ Acknowledgements.
The authors acknowledge helpful conversations with T. Senthil. J.P. acknowledges the support from Florida State University through the FSU Quantum Postdoctoral Fellowship. This research is supported by the National Science Foundation Grant No. DMR-2315954. The portion of the work conducted at Northeastern University was partially supported by NSF CAREER Award No. DMR-2439118. Most of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556, and the State of Florida.
@@ Appendix B Chern-Simons term in the third order SW map
In this Appendix, we show that the diagonal term in the third-order SW map for the gauge potential, $A_{\mu}^{(3,3)}$ in noncommutative phase space can be written as a sum of a total derivative and a term of the Chern-Simons form. For the case of 6+1-D phase-space, we obtain the topological CS term. From the recursive relation for diagonal terms Fidanza , we have
@@
with the understanding that the $\epsilon^{\mu\nu\sigma\rho}$ antisymmetrizes nonzero indices in groups of four. For example, in 6+1-D, it will antisymmetrize the three groups
@@
when the Greek indices take nonzero values $(1,\ldots,6)$. For the first term in Eq. (98), we verified the identity